### v2

@@ 3.1.2 In-Intervention Procedure
Traditional robotic surgery has primarily focused on developing task-specific policies for surgical actions such as suturing [441, 442, 443], tissue manipulation [444, 445], and gauze cutting [446]. The rapid advancements in LLMs and VLAs further enhance the intelligence and versatility of them.
Surgical planner. In conventional clinical settings, meticulous surgical planning is made by surgeons at the initial phase of surgery. With EmAI, this process can be streamlined automatically. Emerging VLA models have been applied in this field to generate actionable surgical plans [447]. They allow robots to understand visual cues and instructions in natural language, enabling more intuitive and flexible task execution [123, 448]. In addition, some VLA models facilitate efficient imitation learning to learn from visual demonstrations to learn complex actions. By interpreting human demonstration videos, EmAI systems are able to recognize fine-grained visual concepts including instruments, verbs, and targets [449] and create actionable task plans [450] for robotic systems in surgical settings.
Moreover, surgical planners powered by Vision-Language Models (VLMs) [451], such as SurgicalGPT [452] and LLaVA-Surg [453], can benefit both practical and educational purposes by offering assistance and evaluation during surgical procedures. Through quantifying surgical performance by decomposing complex surgeries into discrete gestures, the impact of different surgical gestures on patient outcomes can be statistically analyzed and used to predict postoperative results through EmAI systems [454]. Similarly, other studies [455] also support that developing precise, computer-aided surgical planning can help improve patient outcomes. It is suggested that equipping the da Vinci system [456] with AI-based perceptual capabilities can enhance its ability to understand the operating scenarios, plan surgeries, and execute better operations [436].
@@
Surgical operator. Skillful operators are at the core of every surgical procedure, where precision and timely intervention are crucial. Challenges often arise due to the intricate nature of surgeries and the finite availability of expert operators, which can lead to bottlenecks in patient care. To enhance this, robotic surgery has been a promising trend to offer enhanced precision, dexterity, and minimally invasive approaches to common surgical procedures, leading to improved patient outcomes [463]. In healthcare, a dedicated robot is often more practical and reliable, and has been extensively studied in a variety of settings and shown to provide overall clinical benefits [464].
@@
Figure 6: An EmAI system for surgical robotics [465]: it combines strategic planning with precise actuation. This structure optimizes the execution of complex surgical tasks, from detailed pre-operative mappings to real-time adjustments during surgery.
@@
For example, in laparoscopic surgery, there are robotic systems autonomously performing intestinal anastomoses [466] and rectal cancer resection [467] under human oversight. In retinal microsurgery, comprehensive EmAI systems [468, 469] have been developed to integrate real-time surgical object tracking, segmentation, and model predictive control for intraoperative navigation in delicate microsurgical environments wherein precision is decisive.
@@
For general surgical skills, the potential of robotics has also been widely demonstrated. Robots driven by deep learning and active sensing strategies are trained to use needles through the HOUSTON algorithm [470], which can localize, grasp, and hand over unmodified surgical needles to complete precise manipulation. Their good suturing skill that requires a stable and delicate control system has also been proved by the SmartArm robotic system for a neonatal chest surgery [435]. Similarly, for thread detection and interactive perception, real-time robotic approach [437] is presented for autonomously shortening the tail of a surgical suture using a self-supervised learning framework. In complex surgical scenes, multiple EmAI subsystems may collaborate to integrate multi-source perception and support robotic surgery, requiring a robust control system to direct and coordinate their respective perception-cognition-communication-action loops [471]. Beyond visual perception, other modalities including kinematic data [472], audio data [473], language instructions [102], and tactile perception [474] are integrated to form a holistic view of real-world surgical environments.
@@
Intelligent surgeon-robot interface. Intuitive user interfaces and advanced responsive systems can significantly bridge the gap between surgeons and robotic tools, reducing the difficulty of surgeon-robot collaboration and facilitating communication and control as a “translator”. Two kinds of intuitive control and sensory feedback approaches were explored. A novel voice control interface [475] for surgical robots using Whisper speech recognition technology allows surgeons to command robots verbally. Sensory gloves [476] that can translate natural hand movements to control robotic tools strive to provide the tactile feedback that is crucial during surgery but is often missing in robotic operations. Furthermore, human-in-the-loop learning systems [477, 478] involve human interaction to guide and refine the learning process of AI models. These approaches ensure that EmAI is continually improved by incorporating human expertise and feedback for better decision making, which is particularly effective and intuitive in the knowledge-intensive surgical environment. An interactive system that allows surgeons to guide the robot’s learning process during surgery can be more reliable by reducing the error rate, while at the same time being less fully automated.
@@
There are two core components of a surgical navigator, one is the ability to understand and contextualize visual scenes, and the other is safety-oriented path planning and obstacle avoidance. Surgical environments are crowded, and tasks require high precision, making it crucial for AI to correctly identify anatomical landmarks, surgical tools, and other visual cues. Advanced VLN models combine vision transformers and LLMs to enable contextual understanding from both visual and textual inputs [123]. In addition, navigation systems must prioritize safety above all in a surgical setting. AI-driven robots need to navigate within confined spaces while avoiding obstacles, such as surgical instruments, medical staff, or sensitive patient tissues. Path-planning algorithms tailored to the surgical domain incorporate safety constraints and predictive models that anticipate potential obstructions [483, 484]. Techniques such as depth sensing and 3D scene reconstruction are employed to enhance spatial awareness [485, 486, 487], which allows robots to navigate autonomously while ensuring they remain within a safe operating range.
@@
Real-time operation consultant. The modern surgical environment typically involves processing vast amounts of real-time information. Complex procedures generate extensive visual and contextual data, such as live video feeds, laparoscopic images, and robotic surgery footage, which must be rapidly interpreted under high-pressure conditions. Simultaneously, even experienced surgeons may encounter scenarios that exceed their expertise, and reliance on human judgment alone can lead to inefficiencies or errors. To bridge these gaps, AI-powered question-answering approach that are capable of interpreting surgical contexts and providing precise and in-time answers is an essential innovation. Some EmAI systems capable of answering questions based on visual data from surgical environments have emerged [488, 489, 490]. These functionalities are typically enabled by VQA and image captioning approaches, and recent advancements [453, 491, 448, 492, 493, 494, 495] employing novel Vision-Language Pretraining (VLP) techniques in surgery-specific VQA and image captioning tasks have further enhanced the capabilities of EmAI systems, where pre-trained multimodal models are fine-tuned with surgery-specific VQA datasets.
@@
Surgical operation coach. The increasing complexity of surgical procedures, combined with the limited availability of experienced mentors for young trainees, has created significant challenges in providing adequate education for novice surgeons. This training gap exacerbates the global shortage of skilled surgical professionals. Robotic surgical operation coaches, integrated with advanced EmAI algorithms, offer a promising solution to this pressing issue. Researchers have developed EmAI models capable of real-time recognition and prediction of surgical gestures and trajectories [496, 497, 498, 499, 500]. To achieve a comprehensive understanding of the surgical environment, it is crucial to undertake a multi-granular analysis of surgical activities. This includes long-term tasks, such as recognizing surgical phases and steps, as well as short-term tasks, such as segmenting surgical instruments and detecting atomic visual actions [501]. Through these advancements, surgical operation coaches can analyze the intricate sequences of surgical activities, providing objective evaluations of surgical skills [502, 503, 504, 505, 506, 507]. These evaluations deliver constructive feedback, enabling trainees to refine their techniques and accelerate their learning curve.
@@
Patient digital twin. A patient digital twin represents a detailed, dynamic model of a patient’s biological systems or a part thereof (e.g., anatomy), created using comprehensive and accurate medical data such as imaging studies, physiological measurements, and diagnostic results [114]. It is often combined with advanced visualization devices (e.g., VR/AR) to provide interactive observations, and surgeons can plan, simulate, and optimize surgical pathways based on them [508, 509]. AI-based digital twin creation technology facilitates the identification of surgical patterns [438], prediction of complications or procedural outcomes [509], and generation of medical reports [510, 511]. This technology is often employed to accelerate the learning curve of novice surgeons, helping them familiarize themselves with anatomical structures, surgical procedural contexts, and disease progression [512, 513, 514, 515, 516]. Furthermore, multiple downstream applications have benefited from the surgery digitization process [517]. Clinicians can practice surgical skills and explore human anatomy without relying on real anatomical models using digitalized surgical platforms [518], overcoming the limitation of high costs and scarce training samples. It also offers a digital and accurate platform for robots to train on, improving safety and reliability before deployment in actual surgical procedures [519]. Another approach involves photorealistic surgical images [520] and videos [521] synthesis to benefit the training process of EmAI systems, mitigating challenges related to the high costs and ethical concerns of obtaining surgical data.
@@
The developments in EmAI systems for operation and related supporting roles are reshaping surgical practices by enhancing efficiency and precision. These systems provide valuable products and analysis, crucial for both current surgical procedures and the training of future surgeons. However, EmAI-based interventions encompass more than just surgical applications, which are crucial; they also extend to interventional fields including mental health therapy.
@@
Mental healer. EmAI-powered systems leverage advanced emotional recognition, speech analysis, and behavioral pattern detection to assess mental health conditions in real time [522, 523]. Virtual therapists, driven by AI, provide personalized cognitive behavioral therapy (CBT), mindfulness training, and emotional counseling, making mental health support more accessible to underserved populations [524, 525, 526, 527, 528]. Beyond digital therapy platforms, EmAI is being integrated into immersive environments, such as VR-based exposure therapy for phobias or PTSD [529, 530, 531, 532, 533], offering dynamic adjustments to therapy based on physiological feedback like heart rate or eye tracking. EmAI also enhances long-term monitoring through wearable technologies, detecting early warning signs of mental health decline and enabling timely interventions [534, 535, 536, 537], creating a seamless ecosystem of continuous, adaptive, and personalized mental health care.
@@ 3.1.3 Post-Intervention Stage
Intelligent exoskeleton. Emerged EmAI systems [539, 540] have been designed to assist patients in regaining mobility and strength during physical rehabilitation, offering guided exercises, precise motion assistance, and adaptive feedback for individuals with mobility impairments. For example, exoskeleton robots help patients with gait training [541, 542], and hand rehabilitation robots assist in restoring hand function [543]. Upper limb rehabilitation robots have been designed to support patients’ arms during daily activities, promoting neuroplasticity and functional recovery [544, 545, 546]. These technologies are particularly beneficial for stroke or hemiplegia survivors and individuals with spinal cord injuries, offering them the potential for increased independence and improved quality of life [547]. These applications are typically developed with fine control approaches and interaction strategies [548, 549].
Customized recovery coach. For balance exercises on trunk rehabilitation robots, the EmAI systems [550, 551] have shown the ability to automatically capture real-time motions and assess patients’ conditions. They adjust controllers and optimize training intensity according to the individual capabilities and needs of each patient, thereby enhancing personalized rehabilitation sessions. In addition, EmAI systems with the ability to continuously learn can evolve from patients’ past behavior and responses, improving their adaptability and responsiveness to patient-specific therapeutic needs [552, 553]. Moreover, EmAI systems can provide and update safe and personalized rehabilitation training regimens. In particular, they can predict critical rehabilitation metrics [554], enable more accurate assessment of patient recovery, and recommend optimized postoperative rehabilitation plans [555, 556] based on real-time monitoring of patient motor performance and physiological feedback [458, 459, 460].
@@
Health monitoring wearable. Smart wearable devices are increasingly integrated with EmAI systems, utilizing MLLMs capability [567] and enabling the real-time tracking of vital health metrics such as heart rate, oxygen saturation, and electrodermal activity [568, 569, 570]. These devices can dynamically adjust to the unique physiological profiles of individual patients, offering tailored health insights and alerts [571]. Through continuous learning algorithms, EmAI systems evolve to better predict patient-specific health events, such as detecting early signs of infection, Parkinson’s disease or cardiovascular issues [572, 573]. Furthermore, these wearables can communicate with healthcare providers to ensure timely interventions, enhancing patient safety and recovery outcomes [574]. By leveraging real-time data and EmAI, these wearables not only monitor but also proactively manage postoperative care, optimizing recovery trajectories.
Cognitive rehabilitation tool. In the realm of postoperative cognitive recovery, EmAI-equipped tools can customize cognitive exercises based on real-time analysis of patients’ performance and progress [575, 574]. By integrating sensors and interactive software, EmAI systems provide a responsive platform that adjusts tasks according to cognitive load and patient capability [576]. The continuous adaptation helps in crafting highly effective rehabilitation sessions that can address specific cognitive deficits more precisely [577]. Additionally, EmAI tools can predict and monitor cognitive recovery trajectories, offering insights that guide further therapy adjustments [578]. Such tools not only support faster cognitive rehabilitation but also ensure it is engaging and aligned with patients’ specific therapeutic needs.

### v3

@@ 3.1.2 In-Intervention Procedure
Traditional robotic surgery has primarily focused on developing task-specific policies for surgical actions such as suturing [441, 442, 443], tissue manipulation [444, 445], and gauze cutting [446]. The rapid advancements in LLMs and VLAs have further enhanced the intelligence and versatility of these systems.
Surgical planner. In conventional clinical settings, meticulous surgical planning is made by surgeons at the initial phase of surgery. With EmAI, this process can be automated and streamlined. Emerging VLA models have been applied in this field to generate actionable surgical plans [447]. They allow robots to understand visual cues and instructions in natural language, enabling more intuitive and flexible task execution [123, 448]. In addition, some VLA models facilitate efficient imitation learning to learn from visual demonstrations to learn complex actions. By interpreting human demonstration videos, EmAI systems are able to recognize fine-grained visual concepts including instruments, verbs, and targets [449] and create actionable task plans [450] for robotic systems in surgical settings.
Moreover, surgical planners powered by Vision-Language Models (VLMs) [451], such as SurgicalGPT [452] and LLaVA-Surg [453], can benefit both practical and educational purposes by offering assistance and evaluation during surgical procedures. Through quantifying surgical performance by decomposing complex surgeries into discrete gestures, the impact of different surgical gestures on patient outcomes can be statistically analyzed and used to predict postoperative results through EmAI systems [454]. Similarly, other studies [455] support the idea that developing precise, computer-aided surgical planning can help improve patient outcomes. It is suggested that equipping the da Vinci system [456] with AI-based perceptual capabilities can enhance its ability to understand the operating scenarios, plan surgeries, and execute better operations [436].
@@
Surgical operator. Skillful operators are at the core of every surgical procedure, where precision and timely intervention are crucial. Challenges often arise due to the intricate nature of surgeries and the finite availability of expert operators, which can lead to bottlenecks in patient care. To enhance this, robotic surgery has been a promising trend to offer enhanced precision, dexterity, and minimally invasive approaches to common surgical procedures, leading to improved patient outcomes [463]. In healthcare, a dedicated robot is often more practical and reliable and has been extensively studied in a variety of settings, demonstrating overall clinical benefits [464].
@@
Figure 6: An EmAI system for surgical robotics: it combines strategic planning with precise actuation. This structure optimizes the execution of complex surgical tasks, from detailed pre-operative mappings to real-time adjustments during surgery. Source: Reproduced from [465]. Copyright © Boston Engineering 2024.
@@
For example, in laparoscopic surgery, there are robotic systems autonomously performing intestinal anastomoses [466] and rectal cancer resection [467] under human oversight. In retinal microsurgery, comprehensive EmAI systems [468, 469] have been developed to integrate real-time surgical object tracking, segmentation, and model predictive control for intraoperative navigation in delicate microsurgical environments where precision is decisive.
@@
For general surgical skills, the potential of robotics has also been widely demonstrated. Robots driven by deep learning and active sensing strategies have been trained to use needles through the HOUSTON algorithm [470], which can localize, grasp, and hand over unmodified surgical needles to complete precise manipulation. Their exceptional suturing skill that requires a stable and delicate control system has also been proved by the SmartArm robotic system for a neonatal chest surgery [435]. Similarly, for thread detection and interactive perception, real-time robotic approach [437] has been presented for autonomously shortening the tail of a surgical suture using a self-supervised learning framework. In complex surgical scenes, multiple EmAI subsystems may collaborate to integrate multi-source perception and support robotic surgery, requiring a robust control system to direct and coordinate their respective perception-cognition-communication-action loops [471]. Beyond visual perception, other modalities including kinematic data [472], audio data [473], language instructions [102], and tactile perception [474] are integrated to form a holistic view of real-world surgical environments.
@@
Intelligent surgeon-robot interface. Intuitive user interfaces and advanced responsive systems can significantly bridge the gap between surgeons and robotic tools, reducing the difficulty of surgeon-robot collaboration and facilitating communication and control as a “translator”. Two kinds of intuitive control and sensory feedback approaches were explored. A novel voice control interface [475] for surgical robots using Whisper speech recognition technology allows surgeons to command robots verbally. Sensory gloves [476] that can translate natural hand movements to control robotic tools aim to provide the tactile feedback that is crucial during surgery but is often missing in robotic operations. Furthermore, human-in-the-loop learning systems [477, 478] involve human interaction to guide and refine the learning process of AI models. These approaches ensure that EmAI is continually improved by incorporating human expertise and feedback for better decision making, which is particularly effective and intuitive in the knowledge-intensive surgical environment. An interactive system that allows surgeons to guide the robot’s learning process during surgery can be more reliable by reducing the error rate, while at the same time being less fully automated.
@@
There are two core components of a surgical navigator, one is the ability to understand and contextualize visual scenes, and the other is safety-oriented path planning and obstacle avoidance. Surgical environments are crowded, and tasks require high precision, making it crucial for AI to correctly identify anatomical landmarks, surgical tools, and other visual cues. Advanced VLN models combine vision transformers and LLMs to enable contextual understanding from both visual and textual inputs [123]. In addition, navigation systems must prioritize safety above all in a surgical setting. AI-driven robots need to navigate within confined spaces while avoiding obstacles, such as surgical instruments, medical staff, or sensitive patient tissues. Path-planning algorithms tailored to the surgical domain incorporate safety constraints and predictive models that anticipate potential obstructions [483, 484]. Techniques such as depth sensing and 3D scene reconstruction are employed to enhance spatial awareness [485, 486, 487], allowing robots to navigate autonomously while ensuring they remain within a safe operating range.
@@
Real-time operation consultant. The modern surgical environment typically involves processing vast amounts of real-time information. Complex procedures generate extensive visual and contextual data, such as live video feeds, laparoscopic images, and robotic surgery footage, which must be rapidly interpreted under high-pressure conditions. Simultaneously, even experienced surgeons may encounter scenarios that exceed their expertise, and reliance on human judgment alone can lead to inefficiencies or errors. To bridge these gaps, AI-powered question-answering approach that is capable of interpreting surgical contexts and providing precise and in-time answers is an essential innovation. Some EmAI systems capable of answering questions based on visual data from surgical environments have emerged [488, 489, 490]. These functionalities are typically enabled by VQA and image captioning approaches, and recent advancements [453, 491, 448, 492, 493, 494, 495] employing novel Vision-Language Pretraining (VLP) techniques in surgery-specific VQA and image captioning tasks have further enhanced the capabilities of EmAI systems, where pre-trained multimodal models are fine-tuned with surgery-specific VQA datasets.
@@
Surgical operation coach. The increasing complexity of surgical procedures, combined with the limited availability of experienced mentors for young trainees, has created significant challenges in providing adequate education for novice surgeons. This training gap further exacerbates the global shortage of skilled surgical professionals. Robotic surgical operation coaches, integrated with advanced EmAI algorithms, offer a promising solution to this pressing issue. Researchers have developed EmAI models capable of real-time recognition and prediction of surgical gestures and trajectories [496, 497, 498, 499, 500]. To achieve a comprehensive understanding of the surgical environment, it is crucial to undertake a multi-granular analysis of surgical activities. This includes long-term tasks, such as recognizing surgical phases and steps, as well as short-term tasks, such as segmenting surgical instruments and detecting atomic visual actions [501]. Through these advancements, surgical operation coaches can analyze the intricate sequences of surgical activities, providing objective evaluations of surgical skills [502, 503, 504, 505, 506, 507]. These evaluations provide constructive feedback, enabling trainees to refine their techniques and accelerate their learning curve.
@@
Patient digital twin. A patient digital twin represents a detailed, dynamic model of a patient’s biological systems or a part thereof (e.g., anatomy), created using comprehensive and accurate medical data such as imaging studies, physiological measurements, and diagnostic results [114]. It is often combined with advanced visualization devices (e.g., VR/AR) to provide interactive observations, allowing surgeons to plan, simulate, and optimize surgical pathways based on them [508, 509]. AI-based digital twin creation technology facilitates the identification of surgical patterns [438], prediction of complications or procedural outcomes [509], and generation of medical reports [510, 511]. This technology is often employed to accelerate the learning curve of novice surgeons, helping them familiarize themselves with anatomical structures, surgical procedural contexts, and disease progression [512, 513, 514, 515, 516]. Furthermore, multiple downstream applications have benefited from the surgery digitization process [517]. Clinicians can practice surgical skills and explore human anatomy without relying on real anatomical models using digitalized surgical platforms [518], overcoming the limitation of high costs and scarce training samples. It also offers a digital and accurate platform for robots to train on, improving safety and reliability before deployment in actual surgical procedures [519]. Another approach involves photorealistic surgical images [520] and videos [521] synthesis to benefit the training process of EmAI systems, mitigating challenges related to the high costs and ethical concerns of obtaining surgical data.
@@
The developments in EmAI systems for operation and related supporting roles are reshaping surgical practices by enhancing efficiency and precision. These systems provide valuable products and analysis, crucial for both current surgical procedures and the training of future surgeons. However, EmAI-based interventions encompass more than just crucial surgical applications; they also extend to interventional fields such as mental health therapy.
@@
Mental healer. EmAI-based systems leverage advanced emotional recognition, speech analysis, and behavioral pattern detection to assess mental health conditions in real time [522, 523]. Virtual therapists, driven by AI, provide personalized cognitive behavioral therapy (CBT), mindfulness training, and emotional counseling, making mental health support more accessible to underserved populations [524, 525, 526, 527, 528]. Beyond digital therapy platforms, EmAI is being integrated into immersive environments, such as VR-based exposure therapy for phobias or PTSD [529, 530, 531, 532, 533], providing dynamic adjustments to therapy based on physiological feedback like heart rate or eye tracking. EmAI also enhances long-term monitoring through wearable technologies, detecting early warning signs of mental health decline and enabling timely interventions [534, 535, 536, 537], creating a seamless ecosystem of continuous, adaptive, and personalized mental health care.
@@ 3.1.3 Post-Intervention Stage
Intelligent exoskeleton. Emerging EmAI systems [539, 540] have been designed to assist patients in regaining mobility and strength during physical rehabilitation, offering guided exercises, precise motion assistance, and adaptive feedback for individuals with mobility impairments. For example, exoskeleton robots help patients with gait training [541, 542], and hand rehabilitation robots assist in restoring hand function [543]. Upper limb rehabilitation robots have been designed to support patients’ arms during daily activities, promoting neuroplasticity and functional recovery [544, 545, 546]. These technologies are particularly beneficial for stroke or hemiplegia survivors and individuals with spinal cord injuries, offering them the potential for increased independence and improved quality of life [547]. These applications are typically developed with fine control approaches and interaction strategies [548, 549].
Customized recovery coach. For balance exercises on trunk rehabilitation robots, EmAI systems [550, 551] have shown the ability to automatically capture real-time motions and assess patients’ conditions. They adjust controllers and optimize training intensity according to the individual capabilities and needs of each patient, thereby enhancing personalized rehabilitation sessions. In addition, EmAI systems with the ability to continuously learn can evolve from patients’ past behavior and responses, improving their adaptability and responsiveness to patient-specific therapeutic needs [552, 553]. Moreover, EmAI systems can provide and update safe and personalized rehabilitation training regimens. In particular, they can predict critical rehabilitation metrics [554], enable more accurate assessment of patient recovery, and recommend optimized postoperative rehabilitation plans [555, 556] based on real-time monitoring of patient motor performance and physiological feedback [458, 459, 460].
@@
Health monitoring wearable. Smart wearable devices are increasingly integrated with EmAI systems, utilizing MLLMs capabilities [567] and enabling the real-time tracking of vital health metrics such as heart rate, oxygen saturation, and electrodermal activity [568, 569, 570]. These devices can dynamically adjust to the unique physiological profiles of individual patients, offering tailored health insights and alerts [571]. Through continuous learning algorithms, EmAI systems evolve to better predict patient-specific health events, such as detecting early signs of infection, Parkinson’s disease, or cardiovascular issues [572, 573]. Furthermore, these wearables can communicate with healthcare providers to ensure timely interventions, enhancing patient safety and recovery outcomes [574]. By leveraging real-time data and EmAI, these wearables not only monitor but also proactively manage postoperative care, optimizing recovery trajectories.
Cognitive rehabilitation tool. In the realm of postoperative cognitive recovery, EmAI-equipped tools can customize cognitive exercises based on real-time analysis of patients’ performance and progress [575, 574]. By integrating sensors and interactive software, EmAI systems provide a responsive platform that adjusts tasks according to cognitive load and patient capability [576]. The continuous adaptation helps create highly effective rehabilitation sessions that can address specific cognitive deficits more precisely [577]. Additionally, EmAI tools can predict and monitor cognitive recovery trajectories, offering insights that guide further therapy adjustments [578]. Such tools not only support faster cognitive rehabilitation but also ensure it is engaging and aligned with patients’ specific therapeutic needs.